# Homological shifts of a complementary edge ideal

### v1

@@ 4.2. Trees
Set $\sigma_{0}:=0$ and $\sigma_{k}:=\sum\limits_{j=1}^{k}a_{j}$ for all $k\in[n]$.
We now construct a caterpillar tree $T$ such that $K_{d}(T)=I$. Let $P$ be a path of length $n$ with $n+1$ vertices, labeled in order by $\sigma_{1}+1,\sigma_{2}+1,\ldots,\sigma_{n}+1,\sigma_{n}+2$. We then add leaf edges according to the sequence $(a_{1},\ldots,a_{n})$: attach $\sigma_{1}$ leaf edges to the vertex $\sigma_{1}+1$, whose leaf vertices are $1,\ldots,\sigma_{1}$; for each $2\leq i\leq n$, attach $\sigma_{i}-\sigma_{i-1}-1=a_{i}-1$ leaf edges to the vertex $\sigma_{i}+1$, with leaf vertices $\sigma_{i-1}+2,\ldots,\sigma_{i-1}+a_{i}=\sigma_{i}$.
Denote the resulting graph by $T$. By construction, $T$ is a caterpillar tree with vertex set $\sigma_{n}+2]$. Let $\phi:[\sigma_{n}+1]\rightarrow[\sigma_{n}+2]$ be the map where $\phi(k)$ is the unique vertex $j$ satisfying $j>k$ and $\{j,k\}\in E(T)$. For the role of $\phi$, see Theorem 4.3.
@@
Then, for any $k\in[\sigma_{n}+2]$, the set $\phi^{-1}(k)$ is nonempty if and only if $k=\sigma_{j}+1$ for some $1\leq j\leq n$. Furthermore, for each $j\in[\sigma+1]$, we have $\phi^{-1}(\sigma_{j}+1)=\{\sigma_{j-1}+1,\sigma_{j-1}+2,\ldots,\sigma_{j}\}$. In particular, $|\phi^{-1}(\sigma_{j}+1)|=a_{j}$ for $j\in[n]$. By the proof of Proposition 4.5, $K_{d}(T)=I=I_{\mathbf{a},d}$. Moreover,
$\mathrm{HS}_{|\mathbf{a}|-d}(I_{c}(T)^{|\mathbf{a}|-d})=J_{|\mathbf{a}|-d}(T)=uK_{d}(T)=uI$
for some monomial $u$, as required.    ∎
@@ 4.3. Cycles
Let $G$ be an $n$-cycle with vertices labeled $1,2,\ldots,n$ in cyclic order. For $j=1,\ldots,n-1$, let ${\bf e}_{j}$ denote the edge $x_{j}x_{j+1}$, and set ${\bf e}_{n}=x_{n}x_{1}$. For notational convenience, we also use ${\bf e}_{0}$ to denote ${\bf e}_{n}$ and $x_{0}$ to denote $x_{n}$.
We first consider the case when $G$ is an even cycle. Note that the maximal element (set) of the sets $\operatorname{set}(\beta)$, where $\beta\in\mathcal{G}(I^{s})$ for some $s\geq 1$ is equal to $\{1,2,\ldots,n-2\}$.
@@
Let $G$ be the even cycle of size $n=2m$ and let $I$ be the complementary edge ideal of $G$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$, where $u_{j}\in I(G)$ for all $j$. Then
@@
(1) $\Rightarrow$ Suppose that $2\ell$ belongs to $\operatorname{set}(\beta)$ and $u_{j}\neq x_{2\ell}x_{2\ell+1}$ for any $2\leq j\leq s$. Then there exists $k\geq 2\ell$ such that $2\ell-1$ is even connected to $k$ with respect to $u_{2}\cdots u_{s}$ by the following walk:
$2\ell-1=y_{0}--y_{1}--y_{2}--\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=k.$
Assume further that this walk cannot be shotten. We use induction on $\ell$ to show that $\{x_{1}x_{2},x_{3}x_{4},\ldots,x_{2\ell-3}x_{2\ell-2}\}\subseteq\{u_{2},\ldots,u_{s}\}.$ If $\ell=1$, there is nothing to prove, since $\{x_{1}x_{2},x_{3}x_{4},\ldots,x_{2\ell-3}x_{2\ell-2}\}$ is an empty set. Suppose now that $\ell>1$ It is clear that $y_{1}\in\{2\ell,2\ell-2\}$. But if $y_{1}=2\ell$, then $y_{2}$ has to be $2\ell-1$, for otherwise $u_{j}=y_{1}y_{2}=x_{2\ell}x_{2\ell+1}$ for some $j\geq 2$. This implies the above-mentioned walk can be shorten, a contradiction. Hence $y_{1}=2\ell-2$. From this, it follows that $y_{2}=2\ell-3$. Hence, $y_{1}y_{2}\in\{u_{2},\ldots,u_{s}\}$. We may assume $u_{2}=x_{2\ell-3}x_{2\ell-2}$ without loss of generality. Note that $2\ell-3=y_{2q+2}--y_{2q+3}--y_{2q+4}---\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=k$ is a walk by which $2\ell-3$ is even connected to $k$ with respect to $u_{3}u_{4}\cdots u_{s}$. Note that $2\ell-1$ does not belongs to this walk. Hence, by induction, we conclude that $\{x_{1}x_{2},x_{3}x_{4},\ldots,x_{2\ell-5}x_{2\ell-4}\}\subseteq\{u_{3},\ldots,u_{s}\}$, completing the proof.
$\Leftarrow$ Assume that $u_{1}={\bf e}_{2\ell-1}=x_{2\ell-1}x_{2\ell}$. If $u_{j}=x_{2\ell}x_{2\ell+1}$ for some $2\leq j\leq s$, then, since $2\ell+2\leq 2m$ is a vertex of $G$ such that $2\ell+2$ is adjacent to $2\ell+1$ and $2\ell+2>2\ell$, we conclude that $2\ell\in\operatorname{set}(\beta)$ by Lemma 3.3. If $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-3}\}\subseteq\{u_{2},\ldots,u_{s}\},$ then $2m$ is even connected to $2\ell-1$ with respect to ${\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2\ell-3}$ by the following walk
$2m--\underbrace{1--2}-3--4--\cdots--\underbrace{(2\ell-3)--(2\ell-2)}--(2\ell-1).$
Hence, $2\ell$ belongs to $\operatorname{set}(\beta)$.
The proof of (2) is similar to the proof of (1), we omit it.    ∎
Next, we consider the case when $n$ is odd. Note that there exists $\beta\in\mathcal{G}(I^{s})$ for some $s\geq 1$ such that $\operatorname{set}(\beta)$ is equal to $\{1,2,\ldots,n-1\}$.
@@
Let $G$ be the odd cycle of size $n=2m+1$ and let $I$ be the complementary edge ideal of $G$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$, where $u_{j}\in I(G)$ for all $j$. Then the following conclusions hold:
@@
The proofs of (1) and (2) are similar to that of Lemma 4.7, so we only prove (3).
(3) Suppose that $\{{\bf e}_{1},{\bf e}_{3}\ldots,{\bf e}_{2m-3},{\bf e}_{2m-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$. We may assume that $u_{1}={\bf e}_{2m-1}$ without loss of generality. Then, since $2m+1$ is even connected to 2m12m-1 with respect to $u_{2}\cdots u_{s}$ by the following walk:
$(2m+1)--1--2--3--4--\cdots--(2m-3)--(2m-2)--(2m-1),$
@@
Conversely, if $2m\in\operatorname{set}(\beta)$, we may assume that $u_{1}\in\{x_{2m}x_{2m+1},x_{2m-1}x_{2m}\}$. We need to consider the following cases:
(i) $u_{1}=x_{2m}x_{2m+1}$: By Lemma 3.3, $2m+1$ is even connected to $2m+1$ with respect to $u_{2}u_{3}\cdots u_{s}$. We may assume the following walk is a shortest walk by which $2m+1$ even connected to $2m+1$ with respect to $u_{2}u_{3}\cdots u_{s}$:
$2m+1=y_{0}--y_{1}--y_{2}--\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=2m+1,({\dagger})$
It is clear that $y_{1}\in\{1,2m\}$. We consider the case when $y_{1}=1$ only, since the proof of the other case is similar. Since $y_{1}=1$, we have $y_{2}\in\{2,2m+1\}$. But if $y_{2}=2m+1$, then $2m+1--y_{3}-y_{4}----\cdots--(2m-1)--(2m)--(2m+1)$ is also a walk by which $2m+1$ is even connected to $2m+1$. This is contradicted to the assumption that the walk ($\dagger$}) is shortest. Thus $y_{2}=2$. This implies $y_{3}=3$ or $y_{3}=1$. But if $y_{3}=1$, the walk ($\dagger$}) can be shortened, a contradiction again. It follows that $y_{3}=3$. By analogy, we conclude that $y_{4}=4,y_{5}=5,\cdots,$. Finally, we have $t=m$ and $y_{2m}=2m$. This implies that $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$. In particular, $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{1},u_{2},u_{3},\ldots,u_{s}\}$, as desired.
Remark: (1) In view of the proof of (i), we can conclude that $({\dagger})$ is actually the unique shortest walk through which vertex $2m+1$ is even connected to itself with respect to any product of edges. Moreover, any shortest walk that even connects a vertex in the cycle to itself should have the same structure.
(2) In the case when $G$ is an even cycle, we may prove that there are no walks by which a vertex is even connected to itself with respect to any product of edges, using a similar argument as above. This explains why $\max\{\mbox{pd}(I_{c}(G)^{s}):\ s\geq 1\}=|V(G)|-2$, rather than |V(G)|1|V(G)|-1.
(ii) $u_{1}=x_{2m-1}x_{2m}={\bf e}_{2m-1}$: By Lemma 3.3, $2m+1$ is even connected to 2m12m-1 with respect to $u_{2}u_{3}\cdots u_{s}$. We may assume the following walk is a shortest walk by which $2m+1$ is even connected to 2m12m-1 with respect to $u_{2}u_{3}\cdots u_{s}$:
$2m+1=y_{0}--y_{1}--y_{2}--\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=2m-1.$
If $y_{1}=2m$, then $y_{2}=2m-1$. It follows that $2m-1--y_{3}--y_{4}--y_{5}--y_{6}--\cdots---y_{2t-1}--y_{2t}--y_{2t+1}=2m-1$ is a shortest walk that even connects 2m12m-1 to 2m12m-1 with respect to $u_{2}\cdots u_{2}$. This walk must have the same structure as the walk that even connects $2m+1$ to $2m+1$ discussed in Case (i). Hence, it is $(2m-1)--(2m)--(2m+1)--1--2--3--4--\cdots--(2m-3)--(2m-2)--(2m-1)$, and in particular, $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$.
It remains to consider the case when $y_{1}=1$. In this case, we can conclude, as in case (i), that $y_{2}=2,y_{3}=3,\cdots,$, and that t=m1t=m-1, $y_{2m}=2m-2$. Hence, we have $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$ again.    ∎
@@
Let $G$ be the cycle of size $n$, whose vertices are labelled $1,2,\ldots,n$ successively. Let ${\bf e}_{j}$ denote the edge $x_{j}x_{j+1}$ for $j=1,\ldots,n-1$ and set ${\bf e}_{0}={\bf e}_{n}=x_{n}x_{1}$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$. Let $T=\{u_{1},\ldots,u_{s}\}$. Define
$N_{1}(\beta)=\left\{\begin{array}[]{ll}\max\{[2\ell]:\ \{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq T,1\leq\ell\leq m-1\},&\hbox{ $n=2m$;}\\ \max\{[2\ell]:\ \{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq T,1\leq\ell\leq m\},&\hbox{ $n=2m+1$.}\end{array}\right$.
$N_{2}(\beta)=\left\{\begin{array}[]{ll}\max\{[2\ell+1]:\ \{{\bf e}_{0},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq T,\ 1\leq\ell\leq m-2\},&\hbox{$n=2m$;}\\[2.0pt] \max\{[2\ell+1]:\ \{{\bf e}_{0},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq T,\ 1\leq\ell\leq m-1\},&\hbox{$n=2m+1$.}\end{array}\right$.
By convention, if ${\bf e}_{1}\notin\{u_{1},\ldots,u_{s}\}$ then $N_{1}(\beta)=\emptyset$ and if ${\bf e}_{0}\notin\{u_{1},\ldots,u_{s}\}$ then $N_{2}(\beta)=\emptyset$. Then
$\operatorname{set}(\beta)=N_{1}(\beta)\cup N_{2}(\beta)\cup\left(\{\min{u_{j}}:\ 1\leq j\leq s\}\setminus\{n-1\}\right).$
Before the proof, we made some remarks. If $n=2m+1$ is an odd number, any $\beta\in\mathcal{G}(I^{s})$ is expressed as $\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}$ uniquely. Hence, there is no ambiguity in the definitions of $N_{1}(\beta)$ and $N_{2}(\beta)$. However, if $n$ is even, we should note that the elements of $N_{1}(\beta)$ and $N_{2}(\beta)$ depend on the way $\beta$ is written as $\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}$. For example, let $\beta_{1}=\frac{\alpha^{m}}{{\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2m-1}}$ and $\beta_{2}=\frac{\alpha^{m}}{{\bf e}_{2}{\bf e}_{4}\cdots{\bf e}_{2m}}$. Then $\beta_{1}=\beta_{2}$. However, $N_{1}(\beta_{1})=[2m-2]$, $N_{1}(\beta_{2})=\emptyset$, $N_{2}(\beta_{1})=\emptyset$ and $N_{2}(\beta_{2})=[2m-3]$. Of course, $\operatorname{set}(\beta)$ is independent of the expression of $\beta$. In the above example, note that $\min{\bf e}_{2m-2}=2m-2$. we have $\operatorname{set}(\beta_{1})=\operatorname{set}(\beta_{2})=[2m-2]$. Since the only real relation that occurs among ${\bf e}_{1},\ldots,{\bf e}_{n}$ is ${\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2m-1}={\bf e}_{2}\cdots{\bf e}_{2m-2}{\bf e}_{2m}$, the set $\operatorname{set}(\beta)$ given in Lemma 4.9 is actually independent of the expression of $\beta$.

### v5

@@ 5.2. Trees
Recall from the proof of Proposition 5.4 that for a tree $G$ with $n$ vertices, we use $K_{i}(G)$ to denote the dual polymatroidal ideal of $\mathrm{HS}_{i}(I_{c}(G)^{i})$ with respect to the $n$-tupe $(i,\ldots,i)$. We now construct a caterpillar tree $T$ on vertex set $\sigma_{n}+2]$ such that $K_{d}(T)=I$.
At first, let $P$ be a path of length $n$ with $n+1$ vertices, labeled in order by $\sigma_{1}+1,\sigma_{2}+1,\ldots,\sigma_{n}+1,\sigma_{n}+2$. We then add leaf edges according to the sequence $(b_{1},\ldots,b_{n})$: attach $\sigma_{1}$ leaf edges to the vertex $\sigma_{1}+1$, whose leaf vertices are $1,\ldots,\sigma_{1}$; for each $2\leq i\leq n$, attach $\sigma_{i}-\sigma_{i-1}-1=b_{i}-1$ leaf edges to the vertex $\sigma_{i}+1$, with leaf vertices $\sigma_{i-1}+2,\ldots,\sigma_{i-1}+b_{i}=\sigma_{i}$.
Denote the resulting graph by $T$. By construction, $T$ is a caterpillar tree with vertex set $\sigma_{n}+2]$, and its vertex labeling satisfies the FM condition. Define $\phi:[\sigma_{n}+1]\rightarrow[\sigma_{n}+2]$ by assigning to each $k\in[\sigma_{n}+1]$ the unique vertex $j\in[\sigma_{n}+2]$ such that $j>k$ and $\{j,k\}\in E(T)$. For the role of $\phi$, see Theorem 5.2.
@@
It is straightforward to verify that for any $k\in[\sigma_{n}+1]$, the preimage $\phi^{-1}(k)$ is nonempty if and only if $k=\sigma_{j}+1$ for some $1\leq j\leq n$. Furthermore, for each $j\in[n]$, we have $\phi^{-1}(\sigma_{j}+1)=\{\sigma_{j-1}+1,\sigma_{j-1}+2,\ldots,\sigma_{j}\}$, whence $|\phi^{-1}(\sigma_{j}+1)|=b_{j}$. By the proof of Proposition 5.4, it follows that $K_{d}(T)=I$. Hence, we conclude
$\mathrm{HS}_{|\mathbf{b}|-d}(I_{c}(T)^{|\mathbf{b}|-d})=J_{|\mathbf{b}|-d}(T)=uK_{d}(T)=uI$
for some monomial $u$, as required.
@@ 5.3. Cycles
Let $G$ be an $n$-cycle with vertices labeled $1,2,\ldots,n$ in a consecutive cyclic order. For $j=1,\ldots,n-1$, let ${\bf e}_{j}$ denote the edge $x_{j}x_{j+1}$, and set ${\bf e}_{n}=x_{n}x_{1}$. For notational convenience, we also use ${\bf e}_{0}$ to denote ${\bf e}_{n}$ and $x_{0}$ to denote $x_{n}$.
We first consider the case when $G$ is an even cycle. By the definition of a bipartite graph, it is easy to see that every closed walk in a bipartite graph has even length.
@@
Let $G$ be the even cycle of length $n=2m$ and let $I$ be the complementary edge ideal of $G$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$, where $u_{j}\in I(G)$ for all $j$. Then $\operatorname{set}(\beta)\subseteq[n-2]$, and the following assertions hold:
@@
To prove $\operatorname{set}(\beta)\subseteq[n-2]$, it suffices to show that $n-1=2m-1\notin\operatorname{set}(\beta)$. To this end, by Lemma 4.3, it suffices to verify the following two claims:
(i) 2m22m-2 is not even-connected to $2m$ via any product of edges;
(ii) $2m$ is not even-connected to itself via any product of edges.
Suppose, for contradiction, that $2m$ is even-connected to itself with respect to some product of edges. This would immediately imply the existence of a closed walk of odd length in $G$. However, since $G$ is bipartite, every closed walk in $G$ has even length, a contradiction.
Next, suppose that 2m22m-2 is even-connected to $2m$ with respect to some product of edges. This connectivity would yield a closed walk of odd length starting and ending at $2m$, which again contradicts the fact that all closed walks in a bipartite graph have even length.
We now prove the assertion (1).
(1) $\Rightarrow$” Suppose that $2\ell\in\mathrm{set}(\beta)$ and $u_{j}\neq x_{2\ell}x_{2\ell+1}$ for all $2\leq j\leq s$. By Lemma 4.3, there exists $k>2\ell$ such that $2\ell-1$ is even-connected to $k$ with respect to $u_{2}\cdots u_{s}$, via the walk
$2\ell-1=y_{0}-y_{1}-y_{2}-\cdots-y_{2t-1}-y_{2t}-y_{2t+1}=k.$
We further assume this walk cannot be shortened. We proceed by induction on $\ell$ to show that $\{\mathbf{e}_{1},\mathbf{e}_{3},\dots,\mathbf{e}_{2\ell-3}\}\subseteq\{u_{2},\dots,u_{s}\}$.
For $\ell=1$, the claim is trivial, since the set $\{\mathbf{e}_{1},\mathbf{e}_{3},\dots,\mathbf{e}_{2\ell-3}\}$ is empty. Now suppose $\ell>1$.
It is clear that $y_{1}\in\{2\ell,2\ell-2\}$. If $y_{1}=2\ell$, then $y_{2}$ must equal $2\ell-1$; otherwise, $u_{j}=x_{y_{1}}x_{y_{2}}=x_{2\ell}x_{2\ell+1}$ for some $j\geq 2$, which would contradict the non-shortenable assumption of the walk. Thus $y_{1}=2\ell-2$, whence $y_{2}=2\ell-3$. It follows that $\mathbf{e}_{2\ell-3}=x_{y_{1}}x_{y_{2}}\in\{u_{2},\dots,u_{s}\}$. Without loss of generality, we may set $u_{2}=\mathbf{e}_{2\ell-3}$.
Consider now the subwalk
$2\ell-3=y_{2}-y_{3}-y_{4}-\cdots-y_{2t+1}=k.$
This subwalk certifies that $2\ell-3$ is even-connected to $k$ with respect to $u_{3}\cdots u_{s}$. Note that $2\ell-1$ does not appear in this subwalk: the original walk is non-shortenable, and $y_{0}=2\ell-1$ occurs only at the starting vertex. The inductive hypothesis then yields $\{\mathbf{e}_{1},\mathbf{e}_{3},\dots,\mathbf{e}_{2\ell-5}\}\subseteq\{u_{3},\dots,u_{s}\}$. Combining this with $u_{2}=\mathbf{e}_{2\ell-3}$, we conclude that $\{\mathbf{e}_{1},\mathbf{e}_{3},\dots,\mathbf{e}_{2\ell-3}\}\subseteq\{u_{2},\dots,u_{s}\}$, completing the induction.
$\Leftarrow$ If $u_{j}=x_{2\ell}x_{2\ell+1}$ for some $2\leq j\leq s$, then, since $2\ell+2\leq 2m$ is a vertex of $G$ such that $2\ell+2$ is adjacent to $2\ell+1$, we conclude that $2\ell\in\operatorname{set}(\beta)$ by Lemma 4.3. If $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-3}\}\subseteq\{u_{2},\ldots,u_{s}\},$ then $2m$ is even-connected to $2\ell-1$ with respect to $u_{2}\cdots u_{s}$ by the following walk
$2m-\underbrace{1-2}_{\mathbf{e}_{1}}-\underbrace{3-4}_{\mathbf{e}_{3}}-\cdots-\underbrace{(2\ell-3)-(2\ell-2)}_{\mathbf{e}_{2\ell-3}}-(2\ell-1).$
Hence, $2\ell$ belongs to $\operatorname{set}(\beta)$ by Lemma 4.3.
The proof of (2) is similar to the proof of (1), we omit it.
In view of this lemma, we see that the maximal set among all sets of the form $\operatorname{set}(\beta)$, where $\beta\in\mathcal{G}(I_{c}(G)^{s})$ for some $s\geq 1$, is [n2][n-2]. Next, we consider the case when $n$ is odd.
@@
Let $G$ be the odd cycle of length $n=2m+1$ and let $I$ be the complementary edge ideal of $G$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$, where $u_{j}\in I(G)$ for all $j$. Then the following conclusions hold:
@@
The proofs of (1) and (2) are similar to those of Lemma 5.6, so we only present the proof of (3).
(3) Suppose that $\{{\bf e}_{1},{\bf e}_{3}\ldots,{\bf e}_{2m-3},{\bf e}_{2m-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$. We may assume that $u_{1}={\bf e}_{2m-1}$ without loss of generality. Then, since $2m+1$ is even-connected to 2m12m-1 with respect to $u_{2}\cdots u_{s}$ by the following walk:
$(2m+1)-1-2-3-4-\cdots-(2m-3)-(2m-2)-(2m-1),$
@@
Conversely, if $2m\in\text{set}(\beta)$, then there exists some $j\in[s]$ such that $u_{j}\in\{x_{2m}x_{2m+1},x_{2m-1}x_{2m}\}$. Without loss of generality, we may assume $j=1$. We analyze the following cases:
(i) $u_{1}=x_{2m}x_{2m+1}={\bf e}_{2m}$: By Lemma 4.3, $2m+1$ is even-connected to itself with respect to $u_{2}u_{3}\cdots u_{s}$. We may assume the following is a shortest walk by which $2m+1$ is even-connected to itself with respect to $u_{2}u_{3}\cdots u_{s}$:
$2m+1=y_{0}-y_{1}-y_{2}-\cdots-y_{2t-1}-y_{2t}-y_{2t+1}=2m+1,({\dagger})$
It is clear that $y_{1}\in\{1,2m\}$. We only consider the case when $y_{1}=1$, as the proof for the other case is dual. Since $y_{1}=1$, we have $y_{2}\in\{2,2m+1\}$. However, if $y_{2}=2m+1$, then the sequence $2m+1-y_{3}-y_{4}-\cdots-(2m-1)-(2m)-(2m+1)$ also forms a walk by which $2m+1$ is even-connected to itself. This contradicts the assumption that walk ($\dagger$}) is shortest. Thus, $y_{2}=2$. This implies $y_{3}=3$ or $y_{3}=1$. If $y_{3}=1$, walk ($\dagger$}) can be shortened, which is another contradiction. Therefore, $y_{3}=3$. By analogy, we conclude $y_{4}=4,y_{5}=5,\ldots$. Finally, we have $t=m$ and $y_{2m}=2m$, which implies $\{\mathbf{e}_{1},\mathbf{e}_{3},\ldots,\mathbf{e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$. In particular, $\{\mathbf{e}_{1},\mathbf{e}_{3},\ldots,\mathbf{e}_{2m-1}\}\subseteq\{u_{1},u_{2},\ldots,u_{s}\}$, as required.
(ii) $u_{1}=x_{2m-1}x_{2m}={\bf e}_{2m-1}$: By Lemma 4.3, $2m+1$ is even-connected to 2m12m-1 with respect to $u_{2}u_{3}\cdots u_{s}$. We may assume the following walk is a shortest walk by which 2m12m-1 is even-connected to $2m+1$ with respect to $u_{2}u_{3}\cdots u_{s}$:
$2m-1=y_{0}-y_{1}-y_{2}-\cdots-y_{2t-1}-y_{2t}-y_{2t+1}=2m+1.$
If $y_{1}=2m$, then $y_{2}=2m+1$. It follows that $2m+1-y_{3}-y_{4}-y_{5}-y_{6}-\cdots-y_{2t-1}-y_{2t}-y_{2t+1}=2m+1$ is a shortest walk that even-connects $2m+1$ to itself with respect to $u_{2}\cdots u_{2}$. By the conclusion of (i), it follows that $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$.
It remains to consider the case when $y_{1}=2m-2$. In this case, we can conclude, as in case (i), that $y_{2}=2m-3,y_{3}=2m-4,\cdots,$ and that $y_{2m-3}=2$ and $y_{2m-2}=1$. Hence, we have $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$ again.
By the preceding lemmas, we establish the following lemma, which plays a crucial role in obtaining the main result of this subsection.
@@
Let $G$ be the cycle of length $n$, whose vertices are labelled $1,2,\ldots,n$ successively. Let ${\bf e}_{j}$ denote the edge $x_{j}x_{j+1}$ for $j=1,\ldots,n-1$ and set ${\bf e}_{0}={\bf e}_{n}=x_{n}x_{1}$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$. Let $T:=\{u_{1},\ldots,u_{s}\}$. We define
$N_{1}(\beta)=\begin{cases}\max\left\{[2\ell]\mid\{\mathbf{e}_{1},\mathbf{e}_{3},\ldots,\mathbf{e}_{2\ell-1}\}\subseteq T,\ 1\leq\ell\leq m-1\right\},&n=2m;\\ \max\left\{[2\ell]\mid\{\mathbf{e}_{1},\mathbf{e}_{3},\ldots,\mathbf{e}_{2\ell-1}\}\subseteq T,\ 1\leq\ell\leq m\right\},&n=2m+1,\end{cases}$
where the maximum is taken with respect to set inclusion. We further define
$N_{2}(\beta)=\begin{cases}\max\left\{[2\ell+1]\mid\{\mathbf{e}_{0},\mathbf{e}_{2},\ldots,\mathbf{e}_{2\ell}\}\subseteq T,\ 1\leq\ell\leq m-2\right\},&n=2m;\\[2.0pt] \max\left\{[2\ell+1]\mid\{\mathbf{e}_{0},\mathbf{e}_{2},\ldots,\mathbf{e}_{2\ell}\}\subseteq T,\ 1\leq\ell\leq m-1\right\},&n=2m+1,\end{cases}$
where the maximum is also taken with respect to set inclusion.
By convention, if ${\bf e}_{1}\notin\{u_{1},\ldots,u_{s}\}$ then $N_{1}(\beta)=\emptyset$, and if ${\bf e}_{0}\notin\{u_{1},\ldots,u_{s}\}$ then $N_{2}(\beta)=\emptyset$. Then
$\operatorname{set}(\beta)=N_{1}(\beta)\cup N_{2}(\beta)\cup\left(\{\min({u_{j}}):\ 1\leq j\leq s\}\setminus\{n-1\}\right).$
Before the proof, we made some remarks. If $n=2m+1$ is an odd number, any $\beta\in\mathcal{G}(I^{s})$ is expressed as $\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}$ uniquely. Hence, there is no ambiguity in the definitions of $N_{1}(\beta)$ and $N_{2}(\beta)$. However, if $n$ is even, we should note that the elements of $N_{1}(\beta)$ and $N_{2}(\beta)$ depend on the way $\beta$ is written as $\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}$. For example, let $\beta_{1}=\frac{\alpha^{m}}{{\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2m-1}}$ and $\beta_{2}=\frac{\alpha^{m}}{{\bf e}_{2}{\bf e}_{4}\cdots{\bf e}_{2m}}$. Then $\beta_{1}=\beta_{2}$. However, $N_{1}(\beta_{1})=[2m-2]$, $N_{1}(\beta_{2})=\emptyset$, $N_{2}(\beta_{1})=\emptyset$ and $N_{2}(\beta_{2})=[2m-3]$. Of course, $\operatorname{set}(\beta)$ is independent of the expression of $\beta$. In the above example, note that $\min({\bf e}_{2m-2})=2m-2$. we have $\operatorname{set}(\beta_{1})=\operatorname{set}(\beta_{2})=[2m-2]$. Since the only real relation that occurs among ${\bf e}_{1},\ldots,{\bf e}_{n}$ is ${\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2m-1}={\bf e}_{2}\cdots{\bf e}_{2m-2}{\bf e}_{2m}$, the set $\operatorname{set}(\beta)$ given in Lemma 5.8 is actually independent of the expression of $\beta$.